On tuesday d dollars worth of merchandise was sold. On Wednesday the
Amount of merchandise sold was $150 less than twice the amount of merchandise sold on Tuesday
Which exppression represents the amount of merchandise sold on Wednesday?
2(d-150)
2(150-d)
150-2d
2d-150

Found 2 solutions by Theo, josgarithmetic:
Answer by Theo(13342)

i believe it's selection 4 (2d - 150)

on tuesday, d dollars worth of merchandise was sold.
on wednesday, 150 less than twice the amount of merchandise sold on tuesday was sold.

that appears to be 2d - 150.
2d means twice the amount of dollars sold on tueday.
2d - 150 means 150 less than twice the amount of dollars sold on tuesday.

solution:

y = 2d - 150
y represents the amount sold on wednesday.
d represents the amount sold on tuesday.

as an example:
assume 500 was sold on tuesday.
y = 2d - 150 becomes:
y = 2 * 500 - 150 = 1000 - 150 = 850.

500 was sold on tuesday,
850 was sold on wednesday.
2 * 500 = 1000 - 150 = 850.

example confirms the formula is good.

2d - 150 is your solution.


Answer by josgarithmetic(39617)

tuesday, d
wednesday, 2d-150
simple literal wording translation

The answer is already in the description.